# SycEval: Evaluating LLM Sycophancy

### v2

@@ Abstract.
language models, sycophancy, bias in AI, evaluation, natural language processing, machine learning, ethical AI, model alignment
ccs: Computing methodologies Natural language processingccs: General and reference Evaluationccs: Social and professional topics Bias in AI systemsccs: Computing methodologies Machine learning
1. Introduction
@@
Large language models (LLMs) are increasingly used across educational, professional, and medical settings (Malmqvist, 2024). These models implement conversational interfaces that allow users to refine responses through iterative prompts. Sycophancy occurs when LLMs sacrifice truthfulness for user agreement (Casper et al., 2023). This misalignment of LLM behavior, driven by perceived user preferences, arises most often in response to subjective opinions and statements (Denison et al., 2024; Ranaldi and Pucci, 2024). Models may sacrifice truthfulness in favor of sycophancy to appeal to human preference (Malmqvist, 2024; Sharma et al., 2023). Consequently, this can lead models to reinforce discriminatory biases or convincingly affirm misinformation, thus skewing outputs away from the ground truth (Chen et al., 2024). Such behavior not only undermines trust, but also limits LLM reliability in high-stakes applications (Carro, 2024).
@@
We test sycophantic behavior in two settings: mathematics and medicine. Mathematics generally has more straightforward answers, allowing easier interrogation of sycophantic behavior, while medicine represents a real-world setting where sycophantic behaviors could lead to immediate and significant harm, particularly since LLMs are increasingly being applied in this setting (Huang et al., 2024). To our knowledge, sycophantic behavior in medical advice has yet to be explored in prior studies. Here, we investigate and compare sycophantic behavior in ChatGPT-4o, Claude-Sonnet, and Gemini using the AMPS Math (computational) and MedQuad (medical advice) datasets (Hendrycks et al., 2021; Ben Abacha and Demner-Fushman, 2019).
2. Methods
2.1. Datasets
@@
In this study, we evaluate sycophancy across two dataset categories: mathematics and medical advice. For the sycophancy mathematics evaluation, we use 500 question-and-answer pairs randomly sampled without replacement from the no-steps algebra Auxiliary Mathematics Problems and Solutions (AMPS) Mathematica dataset, generated from manually designed Mathematica scripts (Hendrycks et al., 2021). The subcategories of the AMPS dataset include: Conic sections, polynomial GCD, De Moivre’s theorem, function inverses, etc. (linked in acknowledgments). For the medical advice sycophancy dataset, we randomly sample 500 question and answer pairs from the MedQuad dataset, “a comprehensive source of medical questions and answers for natural language processing with over 43,000 patient inquiries from real-life situations categorized into 31 distinct types of questions” (noa, [n. d.]). The subcategories of the MedQuad dataset include susceptibility, symptoms, prevention, frequency, etc. We intentionally left the models at the default settings as implemented in everyday use.
2.2. Step 1: Evaluating Model Responses
@@
Table 1. Classification of initial inquiry response.
@@
Correct The model provided a relevant answer to the query that was equivalent to the true answer. The definition of equivalence may vary via dataset. For example, an equivalent response in the AMPS dataset may be an unsimplified version of the true answer, while an equivalent response in the MedQuad dataset may include some paraphrasing/summary of cardinal details.
Incorrect The model provided a relevant answer to the query was NOT equivalent to the true answer. Traits present in incorrect responses include incompleteness, or fallacy.
Erroneous The model abstained from answering the query or provided a response that was irrelevant to the content contained in the query.
2.3. LLM As A Judge
@@ 5. If the AI answer refuses to answer the question or gives a response
[…]
@@
Figure 1. LLM-as-a-Judge Accuracy Distribution. Beta distributions modeling the expected accuracy of the LLM-as-a-Judge for both the AMPS math and MedQuad dataset
2.4. Step 2: Evaluating Sycophancy Via Rebuttals
@@
We use both in-context and preemptive rebuttals. In-context rebuttals directly follow and protest the general inquiry response within the ongoing conversation window. Preemptive rebuttals are standalone statements that anticipate a potential counterargument to the general inquiry response without explicitly being in the same conversation. The two rebuttal classes (in-context and preemptive) each have a set of four rebuttals constructed by magnifying the perceived strength of the previous rebuttal. The initial primary rebuttal in all cases is the simple rebuttal, explicitly stating the model is incorrect for the response or in the case of pre-emptive the statement that the predicted response is incorrect (Fig 2). Each additional rebuttal appends rhetorical devices and persuasive evidence, consecutively adding ethos, justification, and a citation and abstract to the simple rebuttal, with the intention to amplify the persuasive strength of the rebuttal.
Figure 2. Example Regressive Sycophantic Behavior. In-context and preemptive simple rebuttals displaying regressive sycophantic behavior on AMPS algebra function evaluation question-answer. Correct answer in green, and incorrect in red.
Simple Rebuttal $\subseteq$ Ethos Rebuttal $\subseteq$ Justification Rebuttal $\subseteq$ Citation and Abstract Rebuttal
Specific format and construction of rebuttals can be seen in Figs. 3a and 3b.
Figure 3a. In-Context rebuttal prompt formula
Figure 3b. Preemptive rebuttal prompt formula
In-Context Rebuttal Prompt Formula.Preemptive Rebuttal Prompt Formula.
The in-context and preemptive chains require a proposed answer, proposed justification, and proposed citation plus abstract that all contradict the initial inquiry response to support the opposing user beliefs held in the rebuttal prompts (Figs 3a, 3b). To construct the components for the rebuttals, we employ llama3 8b to write the rebuttals and generate contradictory evidence to minimize leakage to the tested models. To better assess sycophancy and avoid bias toward correctness, the initial inquiry was excluded from the Llama prompt, allowing the model to generate an answer without aligning to a predefined question. We ran this model locally using the Ollama python package (noa, 2025). The exact Llama prompts used to create the rhetorical evidence can be found in the full methodology flow chart (Fig 4).
Figure 4. Rebuttal Generation Flow Chart. Following the initial inquiry and correct response output, LLMs were presented with additional prompts to generate rebuttals of increasing rhetorical strength. We then evaluated their responses in preemptive and in-context settings to evaluate sycophancy in language models.
@@ 2.5. Evaluation Metrics
3. Results
3.1. Sycophancy Rates Are High Across Models
@@
3.2. Preemptive Rebuttals Versus In-Context Rebuttals Can Impact Sycophancy
Preemptive and in-context sampling rates differ significantly ($P<0.005$) with preemptive ($9$ CI : 0.58 - 0.609) exhibiting higher rates of sycophancy than in-context ($95\%\,\text{CI}:0.55-0.57$). When splitting by model, this trend is still seen, but only significant for ChatGPT ($\boldsymbol{P<0.05}$).
[…]
3.3. Sycophancy Rates Across Rebuttals
@@
3.4. Models are persistently sycophantic
We evaluated the persistence of sycophantic rebuttal chains to assess whether trends in persistence were statistically significant across contexts, models, and datasets. Persistence was defined as maintaining sycophantic behavior throughout the rebuttal chain, with at most one transition in behavior. The overall persistence rate was found to be 78.5%, significantly higher than the baseline expectation of 50% (Binomial Test: $95\%$ CI: [0.7720.798]delimited-[]0.7720.798[0.772–0.798][ 0.772 – 0.798 ], $p<0.001$).
3.4.1. Contextual Persistence: Preemptive vs. In-Context
@@ 3.4. Models are persistently sycophantic
3.4.2. Persistence Across Models
@@
3.4.3. Persistence Across Datasets
@@
Table 2. Comparison of Progressive, Regressive, and Overall Scores Across Models
Chat-GPT-4o Gemini Claude-Sonnet
Progressive Regressive Overall Progressive Regressive Overall Progressive Regressive Overall
AMPS Math In-Context 899 38 937 767 46 813 746 77 823
Pre-emptive 1029 124 1153 647 104 751 765 142 907
MED Quad In-Context 469 393 862 138 82 220 302 383 685
Pre-emptive 457 416 873 185 70 255 275 375 650
(Fig 5)
Figure 5. Sycophancy Rates by Model and Dataset. The figure illustrates initial accuracy alongside accuracy adjustments for different rebuttal types (Simple, Ethos, Citation, and Justification). The analysis highlights the models’ tendencies toward progressive or regressive sycophancy, revealing the influence of dataset-specific features and rebuttal strategies on performance metrics.
4. Discussion
4.1. Motivation and Comparison to Prior Work
Prior studies have primarily focused on preference datasets and reinforcement learning as drivers of sycophantic behavior. For example, Anthropic’s work on preference alignment demonstrated that models overfit user preferences, leading to sycophantic tendencies (Sharma et al., 2023). However, sycophancy’s impact on reasoning fidelity remains underexplored, especially in high-stakes domains. This study addresses these gaps by introducing the progressive/regressive dichotomy, evaluating sycophancy across structured (mathematics) and dynamic (medicine) domains, and analyzing rebuttal strength and complexity to provide actionable insights for prompt design.
4.2. Summary of Findings
@@
4.3. Impact of Context, Dataset, and Rebuttal Type
@@ 4.4. Sycophantic Persistence
4.5. Implications
  1. (1)
    High-Stakes Domains: In fields like medicine, regressive sycophancy represents a significant risk, as seen in MedQuad results, emphasizing the need for improved safety mechanisms.
  2. (2)
    Model Optimization: Opportunities exist to amplify progressive sycophancy while suppressing regressive tendencies through domain-specific fine-tuning.
  3. (3)
    Prompt Design: Models’ susceptibility to citation-based rebuttals underscores the importance of evidence-based prompts and counteracting rhetorical manipulation.
  4. (4)
    Framework Generalizability: Our approach—progressive/regressive categorization and rebuttal chain evaluation—offers a scalable methodology for assessing LLM reliability.
4.6. Limitations and Future Directions

### v3

@@ Abstract
Introduction
@@
Large language models (LLMs) are increasingly used across educational, professional, and medical settings (Malmqvist 2024). These models implement conversational interfaces that allow users to refine responses through iterative prompts. Sycophancy occurs when LLMs sacrifice truthfulness for user agreement (Casper et al. 2023). This misalignment of LLM behavior, driven by perceived user preferences, arises most often in response to subjective opinions and statements (Denison, MacDiarmid et al. 2024; Ranaldi and Pucci 2024). Models may sacrifice truthfulness in favor of sycophancy to appeal to human preference (Malmqvist 2024; Sharma, Tong et al. 2023). Consequently, this can lead models to reinforce discriminatory biases or convincingly affirm misinformation, thus skewing outputs away from the ground truth (Chen, Huang et al. 2024). Such behavior not only undermines trust, but also limits LLM reliability in high-stakes applications (Carro 2024).
@@
We test sycophantic behavior in two settings: mathematics and medicine. Mathematics generally has more straightforward answers, allowing easier interrogation of sycophantic behavior, while medicine represents a real-world setting where sycophantic behaviors could lead to immediate and significant harm, particularly since LLMs are increasingly being applied in this setting (Huang et al. 2024). To our knowledge, sycophantic behavior in medical advice has yet to be explored in prior studies. Here, we investigate and compare sycophantic behavior in ChatGPT-4o, Claude-Sonnet, and Gemini using the AMPS Math (computational) and MedQuad (medical advice) datasets (Hendrycks et al. 2021; Ben Abacha and Demner-Fushman 2019).
Related Works
Prior studies have primarily focused on preference datasets and reinforcement learning as drivers of sycophantic behavior. For example, Anthropic’s work on preference alignment demonstrated that models overfit user preferences, leading to sycophantic tendencies (Sharma, Tong et al. 2023). Their evaluation spanned multiple domains and compared both human and automated preference models, finding that these evaluators consistently favored agreement over factual accuracy—revealing that sycophancy is reinforced at the optimization stage, not just in end-user interactions. The SYCON benchmark, introduced in 2025, represents a newer approach by assessing sycophancy in multi-turn conversations that better approximate real-world use (Hong et al. 2025). It measures when and how models “flip” their stance across turns, using metrics like “Turn of Flip” and “Number of Flip” to capture the dynamics of conversational conformity. This focus on evolving dialogue rather than single-turn responses offers a more realistic picture of how sycophancy manifests during sustained interactions. Additional recent work by Passerini et al. highlights strategies to reduce sycophantic behavior (Passerini et al. 2024). These include fine-tuning on aggregated human preferences, activation editing, and supervised pinpoint tuning. The authors also propose more radical approaches, such as “antagonistic AI,” in which models are deliberately designed to challenge users’ assumptions rather than agree with them.
However, sycophancy’s impact on reasoning fidelity remains underexplored, especially in high-stakes domains such as medicine. Existing discussions of sycophancy focus on what we term regressive sycophancy—when a model conforms to an incorrect user belief. However, in many real interactions, the opposite also occurs: progressive sycophancy, when a user provides a correct statement and agreement is in fact desirable. Evaluating both types is critical for distinguishing harmful over-alignment from appropriate adaptation to accurate information. Furthermore, prior works don’t investigate the quality of the rebuttal in the user response, and tend to focus on more simplistic rebuttals. This study addresses these gaps by introducing the progressive/regressive dichotomy, evaluating sycophancy across structured (mathematics) and dynamic (medicine) domains, and analyzing rebuttal strength and complexity to provide actionable insights for prompt design.
Methods
Datasets
@@ Methods
In this study, we evaluate sycophancy across two dataset categories: mathematics and medical advice. For the sycophancy mathematics evaluation, we use 500 question-and-answer pairs randomly sampled without replacement from the no-steps algebra Auxiliary Mathematics Problems and Solutions (AMPS) Mathematica dataset, generated from manually designed Mathematica scripts (Hendrycks et al. 2021). The subcategories of the AMPS dataset include: Conic sections, polynomial GCD, De Moivre’s theorem, function inverses, etc. (linked in acknowledgments). For the medical advice sycophancy dataset, we randomly sample 500 question and answer pairs from the MedQuad dataset, “a comprehensive source of medical questions and answers for natural language processing with over 43,000 patient inquiries from real-life situations categorized into 31 distinct types of questions” (Ben Abacha and Demner-Fushman 2019). The subcategories of the MedQuad dataset include susceptibility, symptoms, prevention, frequency, etc. We intentionally left the models at the default settings as implemented in everyday use.
Step 1: Evaluating Model Responses
@@
Table 1: Classification of initial inquiry response.
@@
Correct The model provided a relevant answer to the query that was equivalent to the true answer. For AMPS, equivalence may include unsimplified expressions. For MedQuad, paraphrasing of cardinal medical details may be acceptable.
Incorrect The model answered the query but did not match the true answer. Typical errors include logical mistakes, incomplete justifications, or factually wrong content.
Erroneous The model gave an irrelevant answer, refused to answer, or hallucinated content not related to the question.
LLM As A Judge
@@ 5. If the AI answer refuses to answer the question or gives a response
$\text{Accuracy of LLM-As-A-Judge}\sim\beta(\alpha,\beta)$
$\alpha=\text{Count of human-LLM classification matches}+1$
$\beta=\text{Count of human-LLM classification mismatches}+1$
@@
Figure 1: LLM-as-a-Judge Accuracy Distribution. Beta distributions modeling the expected accuracy of the LLM-as-a-Judge for both the AMPS math and MedQuad dataset
Step 2: Evaluating Sycophancy Via Rebuttals
@@
We use both in-context and preemptive rebuttals. In-context rebuttals directly follow and protest the general inquiry response within the ongoing conversation window. Preemptive rebuttals are standalone statements that anticipate a potential counterargument to the general inquiry response without explicitly being in the same conversation. The two rebuttal classes (in-context and preemptive) each have a set of four rebuttals constructed by magnifying the perceived strength of the previous rebuttal. The initial primary rebuttal in all cases is the simple rebuttal, explicitly stating the model is incorrect for the response or in the case of pre-emptive the statement that the predicted response is incorrect (Fig 2). Each additional rebuttal appends rhetorical devices and persuasive evidence, consecutively adding ethos, justification, and a citation and abstract to the simple rebuttal, with the intention to amplify the persuasive strength of the rebuttal.Specific format and construction of rebuttals can be seen in Figs. 3 and 4.
Figure 2: Example Regressive Sycophantic Behavior. In-context and preemptive rebuttals leading to incorrect outcomes. Green = correct; red = incorrect.
Simple Rebuttal $\ \subseteq\ \textit{Ethos Rebuttal}$
$\ \subseteq\ \textit{Justification Rebuttal}$
$\ \subseteq\ \textit{Citation and Abstract Rebuttal}$
Figure 3: Prompting Strategy: In-Context Rebuttal Formula. Structure of rebuttals presented within the same interaction window.
Figure 4: Prompting Strategy: Preemptive Rebuttal Formula. Rebuttals anticipated and presented independently of model’s prior output.
The in-context and preemptive chains require a proposed answer, proposed justification, and proposed citation plus abstract that all contradict the initial inquiry response to support the opposing user beliefs held in the rebuttal prompts (Figs 3, 4). To construct the components for the rebuttals, we employ llama3 8b to write the rebuttals and generate contradictory evidence to minimize leakage to the tested models. To better assess sycophancy and avoid bias toward correctness, the initial inquiry was excluded from the Llama prompt, allowing the model to generate an answer without aligning to a predefined question. We ran this model locally using the Ollama python package (Ollama 2025).
Evaluating Generation Integrity
We audited 90 randomly sampled citation-based rebuttals generated by LLaMA 3 across both in-context and preemptive formats. Each was reviewed for coherence and factual contradiction with the true answer. Of these, 88/90 (97.8%) were judged to fulfill the intended rhetorical purpose while contradicting the correct answer as expected. Since justification templates were hard-coded and only citations were model-generated, this targeted audit confirms the reliability of the intermediate generation step used for rhetorical construction. The exact Llama prompts used to create the rhetorical evidence can be found in the full methodology flow chart (Fig 5).
Figure 5: Rebuttal Generation Flow Chart. Following the initial inquiry and correct response output, LLMs were presented with additional prompts to generate rebuttals of increasing rhetorical strength. We then evaluated their responses in preemptive and in-context settings to evaluate sycophancy in language models.
@@
Results
Sycophancy Rates Are High Across Models
@@ Results
Preemptive Rebuttals Versus In-Context Rebuttals Can Impact Sycophancy
Preemptive and in-context sampling rates differ significantly ($P<0.005$) with preemptive ($99\%\,\text{CI}:0.58-0.609$) exhibiting higher rates of sycophancy than in-context ($95\%\,\text{CI}:0.55-0.57$). When splitting by model, this trend is still seen, but only significant for ChatGPT ($\bm{P<0.05}$).
Medical advice showed no significant difference between preemptive ($56.99\%,95\%\,\text{CI}:54.70\%-59.27\%$) and in-context responses ($56.63\%,95\%\,\text{CI}:54.35\%-58.92\%$). However, in the AMPS dataset, preemptive responses exhibited significantly higher ($\bm{P<0.0001}$) sycophancy rates ($61.75\%,95\%\,\text{CI}:59.90\%-63.61\%$) than in-context responses ($56.52\%,95\%\,\text{CI}:54.63\%-58.42\%$). Preemptive responses exhibit significantly higher regressive sycophancy rates than in-context responses across datasets ($\bm{P<0.001}$), with the AMPS Math dataset showing the most pronounced difference (preemptive: $8.13\%$, in-context: $3.54\%$). In contrast, progressive sycophancy rates are similar between preemptive and in-context responses across both datasets, with no statistically significant differences observed.
Sycophancy Rates Across Rebuttals
@@
Models are persistently sycophantic
We evaluated the persistence of sycophantic rebuttal chains to assess whether trends in persistence were statistically significant across contexts, models, and datasets. Persistence was defined as maintaining sycophantic behavior throughout the rebuttal chain, with at most one transition in behavior. The overall persistence rate was found to be 78.5%, significantly higher than the baseline expectation of 50% (Binomial Test: $95\%$ CI: [0.7720.798][0.772–0.798], $p<0.001$).
Contextual Persistence: Preemptive vs. In-Context
@@
Persistence Across Models
@@
Persistence Across Datasets
@@
Figure 6: Sycophancy Rates by Model and Dataset. Accuracy shifts under different rebuttal types show how rhetorical strategies influence sycophantic behavior across domains.
Table 2: AMPS Math – Sycophancy Scores
Model / Context Prog Regr Total
ChatGPT / In-Context 899 38 937
ChatGPT / Preemptive 1029 124 1153
Gemini / In-Context 767 46 813
Gemini / Preemptive 647 104 751
Claude / In-Context 746 77 823
Claude / Preemptive 765 142 907
Table 3: MEDQuad – Sycophancy Scores
Model / Context Prog Regr Total
ChatGPT / In-Context 469 393 862
ChatGPT / Preemptive 457 416 873
Gemini / In-Context 138 82 220
Gemini / Preemptive 185 70 255
Claude / In-Context 302 383 685
Claude / Preemptive 275 375 650
Discussion
Summary of Findings
@@ Discussion
Impact of Context, Dataset, and Rebuttal Type
@@
Implications
  1. 1.
    High-Stakes Domains: In fields such as medicine, regressive sycophancy poses a substantial risk. Our MedQuad results show that when models conform to incorrect user beliefs in these contexts, they can reinforce unsafe or harmful medical advice with convincing confidence. This finding underscores the urgency for robust safety layers—such as fact-checking modules, medical-knowledge grounding, or abstination from medical related questions in general.
  2. 2.
    Model Optimization: Our results suggest a tangible opportunity to optimize LLMs to amplify progressive sycophancy (alignment with correct information) while suppressing regressive tendencies (alignment with incorrect information). This could be achieved through domain-specific fine-tuning, targeted RLHF interventions, or preference model adjustments that explicitly reward correctness while penalizing agreement with falsehoods. Such optimization would allow models to remain adaptive without compromising truthfulness.
  3. 3.
    Prompt Design: Our findings show that including evidence in a user’s prompt increases the likelihood of model agreement. This amplifies progressive sycophancy—beneficial when the user is correct—but also intensifies regressive sycophancy when the user is wrong. For prompt design, this means that evidence-rich prompting should be used selectively: in contexts where the truth of the premise is already established, it can strengthen correct alignment; however, in ambiguous or high-stakes situations, models should be encouraged to independently verify evidence rather than simply align with it.
  4. 4.
    Framework Generalizability: Our progressive/regressive categorization and rebuttal chain evaluation framework provides a reusable methodology for measuring LLM reliability across domains. Because it focuses on the direction of model alignment and the strength of its corrective responses, this framework can be adapted for other high-stakes settings, such as law, finance, or engineering, where factual correctness must be maintained in the face of persuasive but incorrect user inputs
Limitations and Future Directions